OpenStudy (anonymous):

Given the function f(x) = 4(x + 1)2 − 3, indicate the shifts that will affect the location of the vertex, and explain what effect they will have. Use complete sentences. f(x−2) f(x) − 2 f(2x) 2•f(x)

OpenStudy (michele_laino):

hint: we can rewrite yuour equation as follows: \[\Large y + 3 = 4{\left( {x + 1} \right)^2}\]

OpenStudy (michele_laino):

where y=f(x)

OpenStudy (michele_laino):

now we perform this traslation: \[\Large \left\{ \begin{gathered} Y = y + 3 \hfill \\ X = x + 1 \hfill \\ \end{gathered} \right.\] where X, ad Y are the new coordinates. after that traslation, we can rewrite your equation as follows: \[\Large Y = 4{X^2}\] Now do you know what is the graph of the function Y= 4 X^2 ?